A company has two manufacturing plants I and II.in a quality control inspection operation,a random sample of 50 items produced by plant I gave 8 defective while a sample of 50 from plant II gave 12 defective construct confidence interval for the difference in the proportion of items produced by the two plants..


1
Expert's answer
2022-01-19T14:20:16-0500

For plant 1,

"n_1=50\\
x_1=8\\
\hat{p_1}={x_1\over n_1}={8\over50}={4\over25}=0.16"

For plant 2,

"n_2=50\\
x_2=12\\
\hat{p_2}={x_2\over n_2}={12\over50}={6\over25}=0.24"

"\alpha=0.05\\
\hat{q_1}=1-\hat{p_1}=1-0.16=0.84\\
\hat{q_2}=1-\hat{p_2}=1-0.24=0.76"


A 95% confidence interval for the difference in  proportion of items is given by,

"C.I=(\hat{p_1}-\hat{p_2})\pm Z_{\alpha\over2}\sqrt{{\hat{p_1}(\hat{q_1})\over n_1}+{p_2(\hat{q_2})\over n_2}}"

Where,

"Z_{\alpha\over2}=Z_{0.05\over2}=Z_{0.025}=1.96"

Therefore,

"C.I=(0.16-0.24)\pm1.96\sqrt{{0.16\times0.84\over50}+{0.24\times0.76\over50}}\\
C.I=-0.08\pm1.96\sqrt{0.006136}\\
C.I=-0.08\pm0.153531943"

Therefore, a 95% confidence interval for the difference in the proportions of items produced by the two plants is,

"(-0.2335, 0.0735)"